User name Yoomin
Log entry time 08:14:59 on December18,2007
Entry number 221792
There was a couple of minute alarm for target temperature and
pressure when we were doing the loop3 Pb target position scan.
BPM x was about +4mm, and it is suspected that the beam was hitting
the frame of the cryo target.
Temperature went up to 20K.
I asked for beam off, and after the beam went off, target became stable at 19K again soon.